# Density-aware Global-Local Attention Network for Point Cloud Segmentation

### v1

@@ IV-C Outdoor Scene Experiment
TABLE IV: Quantitative comparison experimental results of partial segmentation of 3D point cloud dataset ShapeNet-Part [36].
@@
In addition to experiments on public dataset, we collect point cloud data in real scenes for real scenario segmentation experiments. As shown in Figure 6, it is the visualisation result demonstration of the proposed method for segmentation processing on point cloud data of multiple self-produced oil field scenes after pre-training in the outdoor public dataset. Through the visual demonstration, it can be seen that the proposed method can better achieve more equal attention to the large sample categories and small sample categories, and large objects and small objects when processing point clouds. Especially for the long and thin objects such as poles, railings, pipes, etc., which are more distributed in the actual scene of the oil field, they can be identified and processed better.
IV-D Part Segmentation Experiment
Table IV summarises the experimental results for the 3D point cloud partial segmentation task dataset ShapeNet [36], where the comparison methods [21, 15, 16, 2, 5, 6], and the values of the category-averaged intersection and merger ratios versus the instance-averaged intersection and merger ratios of this paper’s method are listed respectively. From this comparison experiment, it can be seen that the proposed method outperforms the other compared methods in a variety of metrics in the partial segmentation task, again thanks to the newly proposed DbGLA attention module. For local areas with different densities, our method uses windows of different sizes in order to compute the local attention within the corresponding local area, and with the addition of the proposed 3D CR-Loss, it further balances the network’s attention to the different parts of each object, thus improving the segmentation accuracy in the partial segmentation task.
IV-E Ablation Experiment
@@
V Conclusions
In this paper, we propose a point cloud segmentation method that fuses density-aware local attention with global attention to build a balanced perception of local detailes and scene context. Using the proposed Category Response Loss, our method achieves a better balance among different categories and sizes of objects, so that it can better discriminate the harder-to-recognise small objects. Quantitative as well as qualitative experimental results on multiple publicly available datasets demonstrate the effectiveness of the proposed attention mechanism and the novel category response loss. The follow-up work can focus on analysing the shortcomings and corresponding improvement methods on researching the existing Transformer-based 3D semantic segmentation network for indoor scenes to the 3D segmentation task of outdoor scenes.

### v2

@@ 4.3 Outdoor Scene Experiment
In addition to experiments on public dataset, we collect point clouds in real scenes for additional segmentation experiments. As shown in Figure 6, it is the qualitative result demonstration of the proposed method for segmentation on point cloud data of multiple self-produced oil field scenes after pre-training in the outdoor public dataset. Through the visual demonstration, it can be seen that the proposed method can better achieve more equal attention to the large sample categories and small sample categories, and large objects and small objects when processing point clouds. Especially for the long and thin objects such as poles, railings, pipes, which are more distributed in the actual scene of the oil field, they can be identified and processed better.
Table 7: Quantitative comparison experimental results of part segmentation of 3D point cloud dataset ShapeNet-Part [49].
@@
Table 8: Ablation studies on indoor point cloud dataset S3DIS [34] Area 5. SLA stands for shifted-window local attention. VGA represent voxel global attention. WCE means weighted cross entropy loss. $\Delta$ represents the improvement in mIoU compared to the method using only shifted-window local attention and weighted cross-entropy loss.
Attention Loss Term mIoU $\Delta$
SLA WCE 70.6 +0.0
SLA WCE and ours CR-Loss 71.0 +0.4
SLA and VGA WCE 71.2 +0.6
SLA and VGA WCE and ours CR-Loss 71.5 +0.9
Ours DbGLA WCE 71.7 +1.1
Ours DbGLA WCE and ours CR-Loss 72.5 +1.9
4.4 Part Segmentation Experiment
Table 7 shows the experimental results for the 3D point cloud part segmentation task dataset ShapeNet [49], where the comparison methods [2, 5, 6, 13, 14, 19], and the Class mIoU and Instance mIoU of the proposed method are listed respectively. From this comparison experiment, it can be seen that the proposed method outperforms the other compared methods in a variety of metrics in the part segmentation task, again thanks to the newly proposed DbGLA attention module. For local areas with different densities, our method uses windows of different sizes in order to compute the local attention within the corresponding local area, and with the addition of the proposed 3D CR-Loss, it further balances the network’s attention to the different parts of each object, thus improving the segmentation accuracy in the part segmentation task.
Table 9: Ablation studies on outdoor point cloud dataset SemanticKITTI [32] validation set. $\Delta$ represents the improvement in mIoU compared to the method using only shifted-window local attention and weighted cross-entropy loss.
Attention Loss Term mIoU $\Delta$
SLA WCE 69.7 +0.0
SLA WCE and ours CR-Loss 70.6 +0.9
SLA and VGA WCE 71.4 +1.7
SLA and VGA WCE and ours CR-Loss 71.9 +2.2
Ours DbGLA WCE 72.2 +2.5
Ours DbGLA WCE and ours CR-Loss 72.9 +3.2
Table 10: Ablation studies of hyperparameters on S3DIS [34] Area 5 and SemanticKITTI [32] validation set.
Value of $\lambda$ 0.1 0.3 0.5 0.7 0.9
S3DIS mIoU 71.3 71.9 72.5 72.2 71.8
SemanticKITTI mIoU 71.8 72.4 72.9 72.5 72.3
Num. of discrete parts 1 3 5 7 9
S3DIS mIoU 71.0 71.7 72.5 72.3 72.1
SemanticKITTI mIoU 70.6 71.5 72.3 72.9 72.8
Table 11: Ablation studies of the number of iterations on S3DIS [34] Area 5 and SemanticKITTI [32] validation set.
Number of iterations 20 40 60 80 100
S3DIS mIoU 35.2 52.8 65.1 70.9 72.5
SemanticKITTI mIoU 18.9 48.3 65.5 70.3 72.9
Number of iterations 110 120 130 140 150
S3DIS mIoU 72.4 72.3 72.2 72.3 72.1
SemanticKITTI mIoU 72.8 72.5 72.7 72.6 72.5
4.5 Ablation Experiment
@@
In addition, Table 10 shows the ablation studies of the two hyperparameters involved in the proposed method. For $\lambda$ which is used to balance the two loss terms, we set $\lambda=\{0.1,0.3,0.5,0.7,0.9\}$ for the experimental group, and it can be seen that the network accuracy reaches its peak when $\lambda=0.5$, so we set the value of this hyperparameter $\lambda$ to $0.5$ by default in our experiments. For the other hyperparameter, the number of discrete parts of the point cloud densities, we set $\{1,3,5,7,9\}$ for the experimental groups. As can be seen from the data in the table, when the value of discrete parts is set to $1$, the network accuracy is the lowest among all the experimental groups, proving the effectiveness of our proposed density-based attention mechanism. Moreover, for the indoor dataset and the outdoor dataset, the network accuracy comes to the peak when the number of discrete parts is set to $5$ and $7$, respectively, and thus we conduct the experiments according to the corresponding hyperparameter settings.
The ablation study on the number of iterations, presented in Table 11, clearly demonstrates that 100 is the optimal value. As shown in the table, the mIoU on both the S3DIS and SemanticKITTI datasets increases steadily with the number of iterations, peaking at 72.5 and 72.9, respectively, at 100 iterations. Crucially, beyond this point, further training yields no improvement; instead, the performance plateaus and even slightly decreases, indicating a signs of overfitting. Therefore, setting the number of iterations to 100 achieves the best performance while ensuring training efficiency.
[…]
5 Conclusions
In this paper, we propose a point cloud segmentation method that fuses density-aware local attention with global attention to build a balanced perception of local details and scene context. Using the proposed Category Response Loss, our method achieves a better balance among different categories and sizes of objects, so that it can better discriminate the harder-to-recognise small objects. Quantitative as well as qualitative experimental results on multiple publicly available datasets demonstrate the effectiveness of the proposed attention mechanism and the novel category response loss. It should be noted that the effectiveness of the method can be influenced by the density and quality of the input point clouds. Furthermore, while the current framework is designed for indoor scenes, its extension to extremely large-scale outdoor environments presents computational and architectural challenges that warrant further investigation. The follow-up work can focus on analysing the shortcomings and corresponding improvement methods on researching the existing Transformer-based 3D semantic segmentation network for indoor scenes to the 3D segmentation task of outdoor scenes.
CRediT authorship contribution statement
Chade Li: Writing - review &\& editing, Writing - original draft, Visualization, Validation, Resources, Methodology, Investigation, Formal analysis. Pengju Zhang: Data curation. Jiaming Zhang: Writing - review &\& editing. Yihong Wu: Supervision.
Acknowledgment
This work was supported by the National Natural Science Foundation of China under Grant No. 62403459, and Beijing Natural Science Foundation under Grant No. L241012.